OpenStudy (anonymous):

use the exact value of sin(75) to find sin (105) by using sin(180-75)

OpenStudy (john_es):

\[\sin(75)=\sin(45+30)=\sin45\cos30+\sin30\cos45=\frac{\sqrt{6}}{4}+\frac{\sqrt{2}}{4}\] So, \[\sin(105)=\sin(180-75)=\sin180\cos75-\sin75\cos180=\sin75=\frac{\sqrt{6}}{4}+\frac{\sqrt{2}}{4}\]

OpenStudy (anonymous):

THANK YOU SOOOOOOOO MUCH... you were a lot of help..

OpenStudy (john_es):

You're welcome ;)